# Eigen-Value: Efficient Domain-Robust Data Valuation via Eigenvalue-Based Approach

### v2

@@ B.1 Dataset
ImageNet. A large-scale image classification dataset with 1,000 classes. For scalability experiments, we use a subset of the training split. Robustness is measured under domain shifts (V2, Sketch, Rendition, Adversarial). For this benchmark, we performed the data valuation experiment using a subset of 30,000 samples.
DomainNet. A large-scale benchmark for multi-domain learning, containing six stylistically distinct domains: clipart, infograph, painting, quickdraw, real, and sketch. We evaluate generalization by holding out one domain as the target and training on the remaining five. For this benchmark, we performed the data valuation experiment using a subset of 2,000 samples for each domain.
@@ B.2 Experiment setting
    Point addition. We sample 2K in distribution examples. We compute values with each method. We form an initial training set of 1K and retrain while adding the highest value samples from the remaining pool. We evaluate on a different target domain.
    Data removal. We sample 1K of 2K in distribution points. We score them, remove the top 50 percent by value, and train on the rest. We evaluate on the target domain. A larger drop in accuracy indicates a better ability to identify low utility samples.
    Instability. We assess sensitivity to small changes in the training set. We fix 290 of 300 indices, resample the remaining 10, repeat valuation five times, and compute the standard deviation of value rankings on the fixed indices.
Baselines and parameters. For KNN Shapley, we use a validation set of 1K examples and set the neighborhood size to 1K. For Data-OOB, we follow the original paper with num models = 800. We train a logistic regression classifier for 30 epochs with a learning rate of 0.01.
@@
Appendix C Supplementary Experiments
C.1 Empirical Validation of the Matching Marginal Assumption
Figure 9: PCA visualization of normalized embeddings sampled (1K each) from different domain sources. The two distributions, corresponding to different domains, partially overlap due to normalization, illustrating that the matching marginal assumption remains applicable in real-world scenarios.
Although the Matching Marginal assumption we used may appear impractical in real-world scenarios, prior work has shown that normalizing embeddings to have zero mean results in the covariance matrices of ID and OOD data sharing identical diagonal elements while differing in off-diagonal elements (Sun et al., Correlation Alignment for Unsupervised Domain Adaptation, in Domain Adaptation in Computer Vision Applications, Springer, 2017). Furthermore, we empirically verified that this condition holds on real-world datasets such as ImageNet and Amazon Reviews, as demonstrated in Figure 9. Specifically, we sampled 1K data points from each domain, computed their normalized embeddings, and projected them into three-dimensional space using PCA. The hyperplanes in the figure represent the PCA subspaces fitted independently to each domain’s embeddings. Despite differences in domain, we observe that normalized embeddings exhibit shared diagonal elements in their covariance matrices, while differing only in their off-diagonal structure. This supports the claim that the assumption employed in our method imposes no critical limitations when applied to real-world data distributions.
C.2 Experiment on Large and Difficult Domain Shift Benchmark

### v3

@@ B.1 Dataset
ImageNet. A large-scale image classification dataset with 1,000 classes. For scalability experiments, we use a subset of the training split. Robustness is measured under domain shifts. For this benchmark, we performed the data valuation experiment using a subset of 30,000 samples.
DomainNet. A large-scale benchmark for multi-domain learning, containing six stylistically distinct domains: clipart (C), infograph (I), painting (P), quickdraw (Q), real (R), and sketch (S). We evaluate generalization by holding out one domain as the target and training on the remaining five. For this benchmark, we performed the data valuation experiment using a subset of 2,000 samples for each domain.
Unlike prior OOD-aware valuation methods that require access to OOD data during estimation [2, 58], EV operates solely on ID data and therefore remains valid even when no OOD samples are available. This makes EV inherently robust to unseen OOD scenarios, a setting that is common in real-world deployments. All experiments use the train split of each dataset. Performance is also evaluated on randomly sampled data from the train split of the target domain (i.e., the domain not used for training), using a fixed random seed of 42 except for the instability ranking experiment.
@@ B.2 Experiment setting
  • Point addition. We sample 2,000 in distribution examples. We compute values with each method. We form an initial training set of 1,000 and retrain while adding the highest value samples from the remaining pool. We evaluate on a different target domain.
  • Data removal. We sample 1,000 of 2,000 in distribution points. We score them, remove the top 50 percent by value, and train on the rest. We evaluate on the target domain. A larger drop in accuracy indicates a better ability to identify low utility samples.
  • Instability. We assess sensitivity to small changes in the training set. We fix 180 of 200 indices, resample the remaining 20, repeat valuation five times, and compute the standard deviation of value rankings on the fixed indices. These five runs use different random seeds (8, 18, 29, 39, 58).
Baselines and parameters. In this work, we limit the EV-integrated methods to LAVA, KNN Shapley, and Data-OOB, selected based on a balance of performance and computational efficiency. For KNN Shapley, we use a validation set of 1,000 examples and set the neighborhood size to 1,000. For Data-OOB, we follow the original paper with num models = 800. We train a logistic regression classifier for 10 epochs with a learning rate of 0.001.
@@ B.3 Weight parameter
C Supplementary Experiments
C.1 Experiment on ImageNet and DomainNet
@@
C.2 Experiment under Indefinite Hessian
When the Hessian is indefinite, Eq. 11 does not directly apply because the minimum eigenvalue can be negative. By following the same derivation as in the positive semi-definite (PSD) case and restricting the analysis to the subspace spanned by negative eigenvectors, a Rayleigh–based upper bound can still be obtained, where the bound depends on the largest negative eigenvalue. For simplicity, we use the same formulation as in the PSD case. Empirically, incorporating EV consistently improved performance even when the Hessian was indefinite. Table 4 reports the results under the same setup as Table 1 using a 2-layer MLP.
Method CIFAR-10 C $\downarrow$
Random 51.9
InfluenceFunction 50.7
Deviation 51.0
LAVA 44.5
KNN Shapley 41.3
Data-OOB 51.8
EV + LAVA 40.1
EV + KNN Shapley 40.2
EV + Data-OOB 44.0
Table 4: Data removal experiment. Comparison of data valuation methods on CIFAR-10 C using a 2-layer MLP model under domain shift. Lower accuracy indicates more effective identification of low-utility samples. The experimental setup, except the model, follows the same configuration as described in Table 1.
C.3 Experiment under Different Embedder
Our experimental pipeline relies on embedding representations, which raises a natural concern that the performance of EV might depend on the quality or choice of the embedder. To examine this, we repeated the CIFAR-10 data-removal experiments (Table 1) using a variety of embedding models. As reported in Table 5, EV consistently maintains strong robustness to OOD data across different embedders. This result indicates that the OOD robustness of EV is not tied to a specific embedding model and that EV remains effective even when alternative feature extractors are used.
Method / Embedder ResNet-18 ResNet-50 ViT-B/16 ViT-L/16 Avg
Random 48.7 47.0 67.0 70.9 58.7
InfluenceFunction 46.3 47.8 67.5 70.2 57.9
Deviation 43.6 46.5 66.3 71.6 57.0
LAVA 42.7 46.3 66.6 69.9 56.4
KNN Shapley 33.8 38.8 60.4 63.3 49.1
Data-OOB 36.3 44.6 67.8 70.4 54.8
EV + LAVA 42.4 43.9 58.1 58.6 50.8
EV + KNN Shapley 33.9 38.6 60.3 62.5 48.8
EV + Data-OOB 33.6 43.6 63.9 66.6 51.9
Table 5: Data removal experiment. Comparison of data valuation methods on CIFAR-10 C using different embedders. Except for the choice of embedder, the experimental setup follows the same configuration as in Table 1.
C.4 Different Removal Rate Results on CIFAR-10 C
While Table 1 reports the results of removing the top 50% highest-value samples, we additionally evaluate a less extreme setting to show that EV is not effective only under large removal rates. Table 6 presents the results when the removal rate is reduced to 10%. EV continues to deliver strong performance at this lower rate, exhibiting consistent improvements across methods. These results indicate that EV is broadly applicable and remains effective under a range of removal rates.
Method CIFAR-10 C $\downarrow$
Random 50.1
InfluenceFunction 49.3
Deviation 50.7
LAVA 50.03
KNN Shapley 48.19
Data-OOB 48.5
EV + LAVA 48.6
EV + KNN Shapley 45.8
EV + Data-OOB 46.9
Table 6: Data removal experiment. Comparison of data valuation methods on CIFAR-10 C using different removal rates (10%). Lower accuracy indicates more effective identification of low-utility samples. The experimental setup, except removal rate, follows the same configuration as described in Table 1.
C.5 Additional Baseline Results on CIFAR-10 C
We extended the data removal experiment of Table 1 by applying the same evaluation procedure to additional baselines on CIFAR-10 C. As summarized in Table 7, we included both DAVINZ [56] and Banzhaf [54] under the identical experimental setup. Across these baselines, incorporating EV led to consistent improvements, indicating that the proposed approach effectively enhances the ability of diverse valuation methods to identify low-utility samples. These results confirm that the benefit of EV is not limited to a specific algorithm but generalizes across methods with different underlying principles.
Method CIFAR-10 C $\downarrow$
Random 47.0
InfluenceFunction 47.8
Deviation 46.5
LAVA 46.3
KNN Shapley 38.8
Data-OOB 44.6
Banzhaf 46.7
DAVINZ 47.3
EV + LAVA 43.9
EV + KNN Shapley 38.6
EV + Data-OOB 43.6
EV + Banzhaf 45.4
EV + DAVINZ 35.6
Table 7: Data removal experiment. Comparison of data valuation methods on CIFAR-10-C under domain shift, including additional results for Banzhaf and DAVINZ methods. The experimental setup follows the same configuration as described in Table 1.
C.6 Additional metrics for the Instability Ranking experiment
In Figure 4, we examine how sensitive each valuation method is to small perturbations in the training subset by computing the standard deviation of the resulting data value rankings across multiple resamples. This analysis illustrates how much the ranking produced by each method fluctuates when the underlying subset used for valuation is slightly altered. Table 8 provides a more comprehensive evaluation based on the rankings obtained in Figure 4. In addition to reporting the standard deviation, we also compute the variance, Spearman correlation, and Kendall correlation across repeated runs. These complementary metrics reveal that EV achieves substantially greater stability than Deviation, demonstrating that EV retains robust OOD performance even when the training subset undergoes minor changes.
Method / Metric Std. $(\downarrow)$ Variance $(\downarrow)$ Spearman $(\uparrow)$ Kendall $(\uparrow)$
Random 48.58 2582.65 0.05 0.03
Deviation 34.25 1489.68 0.45 0.32
EV + LAVA 7.60 75.45 0.97 0.86
EV + KNN Shapley 7.84 81.77 0.97 0.86
EV + Data-OOB 19.18 592.09 0.77 0.62
Table 8: Additional metrics for the Instability Ranking experiment